OpenStudy (anonymous):

Solve for b. \frac{ 1 }{ b-5 }-\frac{ 10 }{ b^2-5b+25 }=\frac{ 1 }{ b+5 }

OpenStudy (anonymous):

\[\frac{ 1 }{ b-5 }-\frac{ 10 }{ b^2-5b+25 }=\frac{ 1 }{ b+5 } \]